OpenStudy (anonymous):

what is 23 base 4 plus 21 base 4?

jimthompson5910 (jim_thompson5910):

1+3 = 4 but 4 doesn't exist in base 4 (much like the number 10 as a single digit doesn't exist in base 10) so this means 1+3 = 10 in base 4

jimthompson5910 (jim_thompson5910):

2+2 = 10 in base 4 as well and we also have a carry over, so we really have 1+2+2 = 1+10 = 11 base 4

jimthompson5910 (jim_thompson5910):

put this all together to get